OpenStudy (anonymous):

Solve for the variable: 2a/3 = 3/4

OpenStudy (anonymous):

a = 9/8

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2a/3 -3/4 = 0 (8a - 9)/12 = 0 a = 9/8

OpenStudy (anonymous):

Simplify, then cross multiply

OpenStudy (anonymous):

I get a = 1/8

OpenStudy (anonymous):

hmm...how? 2(1/8)/3 = 3/4 ??

OpenStudy (anonymous):

ohh... m sorry..:(

OpenStudy (anonymous):

I stand corrected

OpenStudy (anonymous):

It's 9/8. My algebra failed apparently.

OpenStudy (anonymous):

But still, if you cross multiply without simplifying, you will get the same answer